OpenStudy (anonymous):

Find the range of values for p for whih the roots of the roots of the equation x^2 + (2p +1)x = 3 - p^2 are not real. State the range of values for p for which the roots are real.

OpenStudy (anonymous):

in order for evaluate the range of values for p for which roots are real : Discriminant of quadratic must be non-negative

OpenStudy (anonymous):

\[x^2+(2p+1)x+p^2-3=0\]\[\Delta=(2p+1)^2-4\times1\times(p^2-3)=4p+13\]

OpenStudy (anonymous):

and one more thing : in order for evaluate the range of values for p for which roots are not real : Discriminant of quadratic must be negative or \(\Delta<0\)

OpenStudy (anonymous):

oh thanks so much i realised that i made a mistake when i tried to work this out, i put 3-p^2 instead of p^-3!

OpenStudy (anonymous):

oh ok :)